I have a right triangle. One of the legs is 8cm, the other leg is 'x' and the hypotenuse is 'x+4' how do I solve using pythagorean theorem if my hypotenuse is a binomial?
Answer by mathmartian(7)

Since you have a binomial for the hypotenuse, make sure you square the entire binomial.
Pythagorean theorem: a²+b²=c² where a and b are the legs and the c is the hypotenuse
Substitute your legs and hypotenuse into the formula and solve
8²+x²=(x+4)²
64+x²=x²+8x+16
64=8x+16 Subtract x² from both sides
48=8x Subtract 16 from both sides
x=6 Divide both sides by 8
Now you can use x=6 to solve for the leg and hypotenuse.
The leg is 6cm while the hypotenuse is 6+4, or 10cm.